# Constrained Optimization From a Control Perspective via Feedback Linearization

### v2

@@ 6.1 Heterogeneous Logistic Regression
where $\theta$ is the parameter of the regression model. We further define the averaged regression loss $\bar{R}(\theta)$ as
$\textstyle\bar{R}(\theta):=\frac{1}{C}\sum_{c=1}^{C}f_{c}(\theta)$
As suggested in Shen et al. (2022); Hounie et al. (2024), heterogeneity challenges can be addressed by introducing a proximity constraint that links the performance of each individual client, $R_{c}$, to the average loss across all clients, $\bar{R}$. This approach naturally formulates a constrained learning problem:
$\textstyle\min_{\theta}\bar{R}(\theta),~~~~~s.t.~R_{c}(\theta)\!-\!\bar{R}(\theta)\!-\!\epsilon\!\leq\!0,~\forall c\!\in\!\{1,2,\dots,C\}$ (21)
@@
We solve the constrained optimization problem (21) by running the FL-proximal, FL-Newton, and FL-momentum algorithm . Here we set the number of clients to $C=5$ and $|D_{c}|=200$, the data $y_{i}$ is randomly generated from a Bernoulli distribution and $x_{i}$ is generated from a Gaussian distribution whose mean differs among different agents. The results of the numerical simulation are presented in Figure 2. Notably, all algorithms successfully converge to a first order KKT point, with FL-Newton exhibiting the fastest convergence, followed by FL-momentum, which outperforms FL-proximal in terms of convergence speed.
6.2 Optimal Power Flow
where the objective function $f(x)$ represents the power generation cost and the equality constraints $h_{\textup{eq}}(x)$ generally represents the physical law of the power system, i.e., the power flow equations and $h_{\textup{ineq}}$ includes operational limits in terms of voltage, power generation, transmission capacities etc. The optimization variable $x$ generally consists of voltage angles and magnitudes at each bus, and the real and reactive power injections at each generator (see Low (2014) for a detailed introduction on AC OPF).
Figure 3: Result for AC OPF on IEEE-39 bus (left) and IEEE-118 bus (right) bus system
@@ Appendix A Related Literature
As mentioned in the introduction, many papers analyze the performance continuous time optimization algorithms in different settings. For unconstrained cases, various continuous algorithms are studied such as gradient flow Elkabetz and Cohen [2021], Arora et al. [2019], Saxe et al. [2013], Garg and Panagou [2021], Andrei , stochastic gradient Raginsky et al. [2017], Xie et al. [2021], Nesterov acceleration Su et al. [2016], Wilson et al. [2018], Muehlebach and Jordan [2019], and momentum acceleration Polyak [1964], Polyak and Shcherbakov [2017]. Notably, control-theoretic tools, such as integral quadratic constraints (IQC) Lessard et al. [2016], Hu et al. [2017], Proportional Integral Derivative (PID) control Hu and Lessard [2017], are used to bring interesting insights to the convergence property of optimization algorithms. However, for constrained cases, less work is known. One direction focuses on the dynamics of PDGD algorithm, which primarily focuses on the convex setting and establishing global convergence Kose [1956], Qu and Li [2018], Ding and Jovanović [2019], Zeng et al. [2023], Cerone et al. [2024]. However, this approach is primarily limited to convex optimization. For nonconvex optimization, prior research has explored modifications of the gradient flow to accommodate equality constraints (see, e.g., Yamashita [1980], Schropp and Singer [2000], Wang et al. [2003]). The algorithms proposed in these works share similarities with the feedback linearization approach considered in Cerone et al. [2024] and this paper. Notably, Schropp and Singer [2000] also highlighted connections to sequential quadratic programming (SQP). This paper advances the existing literature in several key ways: i) while prior analyses Yamashita [1980], Schropp and Singer [2000], Zhou et al. [2007] primarily focus on equality constraints, our results extend the algorithm to handle inequality constraints; ii) although the algorithms in these works resemble ours, we introduce a novel perspective based on feedback linearization; iii) previous convergence results are largely asymptotic, whereas we establish non-asymptotic convergence rates in terms of the KKT gap; and iv) We establish acceleration results through the momentum method, further contributing to the theoretical understanding of the algorithm.
@@
Note that our convergence results rely on the linear independent constraint qualification (LICQ) assumption (Assumption 1.1). In literature, there are works that seek to replace this restrictive assumption with milder conditions while still maintaining convergence properties, which is generally known as the stabilized SQP (cf. Gill and Robinson [2013], Gill et al. [2017], Hager [1999], Wright [1998], Zhang and Chen [2024]). It is an interesting open question whether we can borrow insights from stabilized SQP to modify our FL algorithms such that the LICQ assumption can be removed.
@@
Augmented Lagrangian Methods (ALMs) (cf. Birgin and Mart´ınez [2014], Curtis et al. [2015]), also known as the method of multipliers, is another popular approach to solve large-scale constrained optimization problems by transforming them into a series of unconstrained problems. This transformation is achieved by augmenting the standard Lagrangian function with a penalty term that penalizes constraint violations, thereby facilitating convergence to the optimal solution.
@@
Appendix B Proof of Theorem 1
@@
Appendix C Proof of Theorem 2 and 3
@@
is the dual problem of (4) (see Lemma 2 in Appendix F). Thus the solution of $\lambda^{\star}$ in (13) serves as the Lagrangian dual variable of (3.2). Then, we have that the $x^{\star}$ is an optimal solution of (4) if and only if
@@
Appendix D Proof of Theorem 4
@@
Appendix E Proof of Theorem 5
@@
E.1 Intuition: optimization with affine constraints
@@ Appendix F Auxiliaries
$\operatorname*{arg\,min}_{x}\nabla f(x_{t})^{\top}(x-x_{t})+\frac{1}{2}(x-x_{t})^{\top}T(x_{t})^{-1}(x-x_{t})+\lambda^{\top}\left(h(x_{t})+J_{h}(x_{t})(x-x_{t})\right)=x_{t}-\eta T(x_{t})(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda),$
@@
Assumption 1.1 along with Assumption 2.2 is a sufficient condition of Assumption 2.1.
@@
From Assumption 1.1 and Assumption 2.2 we have that
@@
Hence, Assumption 2.1 is satisfied by setting $L=\frac{\lambda_{\max}M^{2}+\|\lambda_{\max}(K)h(x(0))\|}{{\lambda_{\min}}}D^{2}$. ∎

### v3

@@ 6 Numerical Verifications
where $\theta$ is the parameter of the regression model. We further define the averaged regression loss $\bar{R}(\theta)$ as $\bar{R}(\theta):=\frac{1}{C}\sum_{c=1}^{C}f_{c}(\theta).$
As suggested in Shen et al. (2022); Hounie et al. (2024), heterogeneity challenges can be addressed by introducing a proximity constraint that links the performance of each individual client, $R_{c}$, to the average loss across all clients, $\bar{R}$. This approach naturally formulates a constrained learning problem  [^3] It is not hard to verify that with probability one Equation (22) satisfies Assumption 2 and 4. Also within a compact region, the function satisfies Assumption 1 (Lipschitz property) as well
$\textstyle\min_{\theta}\bar{R}(\theta),~~~~~s.t.~R_{c}(\theta)-\bar{R}(\theta)-\epsilon\leq 0,~\forall c\in\{1,2,\dots,C\}$ (22)
@@
We solve the constrained optimization problem (22) by running the FL-proximal, FL-Newton, and FL-momentum algorithm . Here we set the number of clients to $C=5$ and $|D_{c}|=200$, the data $y_{i}$ is randomly generated from a Bernoulli distribution and $x_{i}$ is generated from a Gaussian distribution whose mean differs among different agents. The results of the numerical simulation are presented in Figure 2. All algorithms converge to a first-order KKT point. FL-Newton converges fastest owing to its use of second-order (Hessian) information, while among first-order methods, FL-momentum outperforms FL-proximal in convergence speed. More comprehensive numerical results are provided in Appendix B.
@@ 7 Conclusion
Several limitations of our framework remain. First, our analysis ensures convergence of the best-found iterate to a first-order KKT point, but does not distinguish between local optima, global optima, and saddle points. Extending the framework to guarantee convergence to local optima via saddle-point escape mechanisms is an important direction for future work. Second, the applicability of feedback linearization depends on structural assumptions such as the Linear Independence Constraint Qualification (LICQ). Third, while FL-momentum empirically accelerates convergence, our current analysis yields the same $O(1/\sqrt{T})$ rate as its non-accelerated counterpart. Relaxing these assumptions and strengthening convergence guarantees remain promising directions for future research.
Acknowledgement
This work was supported in part by NSF ASCENT under Grant 2328241 and in part by the NSF AI Institute under Grant 2112085. Runyu Zhang was supported by the MIT Postdoctoral Fellowship Program for Engineering Excellence.
@@ Appendix A Related Literature
As mentioned in the introduction, many papers analyze the performance continuous time optimization algorithms in different settings. For unconstrained cases, various continuous algorithms are studied such as gradient flow Elkabetz and Cohen [2021], Arora et al. [2019], Saxe et al. [2013], Garg and Panagou [2021], Andrei , stochastic gradient Raginsky et al. [2017], Xie et al. [2021], Nesterov acceleration Su et al. [2016], Wilson et al. [2018], Muehlebach and Jordan [2019], and momentum acceleration Polyak [1964], Polyak and Shcherbakov [2017]. Notably, control-theoretic tools, such as integral quadratic constraints (IQC) Lessard et al. [2016], Hu et al. [2017], Proportional Integral Derivative (PID) control Hu and Lessard [2017], are used to bring interesting insights to the convergence property of optimization algorithms. However, for constrained cases, less work is known. One direction focuses on the dynamics of PDGD algorithm, which primarily focuses on the convex setting and establishing global convergence Kose [1956], Wang and Elia [2011], Ding and Jovanović [2019], Zeng et al. [2023], Cerone et al. [2025]. However, this approach is primarily limited to convex optimization. For nonconvex optimization, prior research has explored modifications of the gradient flow to accommodate equality constraints (see, e.g., Yamashita [1980], Schropp and Singer [2000], Wang et al. [2003]) as well as inequality constraints Allibhoy and Cortés [2024], Muehlebach and Jordan [2022b, 2023] . The algorithms proposed in these works share similarities with the feedback linearization approach considered in Cerone et al. [2025] and this paper. Notably, Schropp and Singer [2000] also highlighted connections to sequential quadratic programming (SQP). This paper advances the existing literature in several key ways: i) while prior analyses Yamashita [1980], Schropp and Singer [2000], Zhou et al. [2007] primarily focus on equality constraints, our results extend the algorithm to handle inequality constraints; ii) although the algorithms in these works resemble ours, we introduce a novel perspective based on feedback linearization; iii) previous convergence results are largely asymptotic, whereas we establish non-asymptotic convergence rates in terms of the KKT gap; and iv) We establish acceleration results through the momentum method, further contributing to the theoretical understanding of the algorithm. Notably, Muehlebach and Jordan [2023] also considers a momentum-based acceleration for constrained optimization, however, both the continuous-time dynamics and the discretization approach are different, and thus resulting in distinct algorithms.
@@
Note that our convergence results rely on the linear independent constraint qualification (LICQ) assumption (Assumption 3). In literature, there are works that seek to replace this restrictive assumption with milder conditions while still maintaining convergence properties, which is generally known as the stabilized SQP (cf. Gill and Robinson [2013], Gill et al. [2017], Hager [1999], Wright [1998], Zhang and Chen [2024]). It is an interesting open question whether we can borrow insights from stabilized SQP to modify our FL algorithms such that the LICQ assumption can be removed.
@@
Augmented Lagrangian Methods (ALMs) (cf. Birgin and Martínez [2014], Curtis et al. [2015]), also known as the method of multipliers, is another popular approach to solve large-scale constrained optimization problems by transforming them into a series of unconstrained problems. This transformation is achieved by augmenting the standard Lagrangian function with a penalty term that penalizes constraint violations, thereby facilitating convergence to the optimal solution.
@@
Appendix B More Numerical Simulations
B.1 Comparison with other first-order methods
Here we also compare the performance of our algorithm with existing first order methods such as primal dual gradient descent (PDGD) and Augmented Lagrangian method (ALM), as show in the figure below:
Figure 3: Running different constrained optimization algorithm for Problem (22)
From the above figure, we note that our FL-proximal method performs comparably with existing methods such as PDGD and ALM. And as expected, FL-momentum converges faster as it leverages momentum acceleration and FL-Newton performs the best as it leverages the second order information.
B.2 Generalization of the FL algorithm using PI Control
As pointed out by Remark 1, one drawback of the FL-proximal is that it requires calculation of the inverse matrix of $J_{h}(x)J_{h}(x)^{\top}$, which might be inefficient when the number of constraints becomes larger. In reality, there are ways to solve the inverse of $J_{h}(x)J_{h}(x)^{\top}$ approximately using heuristic methods or conjugate gradient.
Assume that instead of calculating $(J_{h}(x)J_{h}(x)^{\top})^{-1}$ we approximate it with some matrix $F(x)\approx(J_{h}(x)J_{h}(x)^{\top})^{-1}$. For example, in the setting where $J_{h}(x)J_{h}(x)^{\top}$ is diagonal dominant, we can heuristically set $F(x)$ to be a diagonal matrix where the diagonal element of $F(x)$ is the inverse of the corresponding entry of $J_{h}(x)J_{h}(x)^{\top}$, i.e.,
$F(x):=\left(\textup{diag}(J_{h}(x)J_{h}(x)^{\top})\right)^{-1}.$ (23)
Thus, it is of great importance to study the algorithm’s robustness towards the approximation error of $F(x)$. We are going to see in this section that it might be beneficial to consider a PI controller for feedback linearization.
We begin by introducing the Approximated FL algorithm for equality constrained optimization as follows:
Approximated FL $\vskip-20.0pt\dot{x}$ $=-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)$ $\lambda$ $=-F(x)(J_{h}(x)T(x)\nabla f(x)-Kh(x))$
Note that the only difference of Approximated FL comparing with FL-proximal is that we replace $(J_{h}(x)J_{h}(x)^{\top})^{-1}$ with its approximation $F(x)$.
As mentioned in Remark 2, apart from considering the feedback linearization with a proportional controller, i.e. $\dot{y}=-Ky$, we can also consider a more general algorithm variant where we also include the integral term and thus formulate the dynamics as
$\textstyle\dot{y}=-K_{p}y-K_{i}\int_{0}^{t}y(s)ds\vskip-3.0pt$
This gives rises to the (Approximated) FL-PI algorithm as follows:
Approximated FL-PI $\vskip-20.0pt\dot{x}$ $=-\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)$ $\lambda$ $=-F(x)\left(J_{h}(x)T(x)\nabla f(x)-\left(K_{p}h(x)+K_{i}\int_{s=0}^{t}h(x(s))ds\right)\right)$
Figure 4: Comparison of Approximated FL and Approximated FL-PI algorithm
To test the robustness of the algorithm towards approximation error, we run both Approximated FL and Approximated FL-PI algorithm on a quadratic programming problem
$\min_{x}\frac{1}{2}x^{\top}Qx+c^{\top}x$
$s.t.~~Ax+b\leq 0$
where $m=10,n=20$, and $A,b,c,Q^{1/2}$ are all random matrices or vectors whose entries are generated following an i.i.d random Gaussian distribution.
Figure 4 demonstrates the algorithm performance for both Approximated FL and Approximated FL-PI. Interestingly, in this setting, Approximated FL fails to converge to the optimal solution, whereas Approximated FL-PI is able to find the optimal solution. This indicates that it is beneficial to include an additional integral controller term into the optimization algorithm.
B.3 More numerical examples: Optimal Power Flow
where the objective function $f(x)$ represents the power generation cost and the equality constraints $h_{\textup{eq}}(x)$ generally represents the physical law of the power system, i.e., the power flow equations and $h_{\textup{ineq}}$ includes operational limits in terms of voltage, power generation, transmission capacities etc. The optimization variable $x$ generally consists of voltage angles and magnitudes at each bus, and the real and reactive power injections at each generator (see Low [2014] for a detailed introduction on AC OPF).
Figure 5: Result for AC OPF on
IEEE-39 bus (left) and IEEE-118 bus (right) bus system
Appendix C Proof of Theorem 1
@@
C.1 Global Convergence for Strongly Convex Settings
Under the assumption that the optimization problem (1) is strongly convex, we are able to establish the result that FL-proximal converges to the global optimal solution with a linear rate. We make the following assumption.
Assumption 7 (Strong Convexity).
The objective function $f(x)$ is strongly convex, i.e. for any $x,y\in\mathbb{R}^{n}$,
$f(x)-f(y)\geq\nabla f(y)^{\top}(x-y)+\frac{\mu}{2}\|x-y\|^{2}.$
The constraint function $h(x)$ in (1) takes the form of $h(x)=Ax+b$ where $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m}$, and $A$ is of full row rank.
Theorem 6.
Under Assumption 7, we have that running FL-proximal ((6) with $T(x)=I$) will give
$f(x)-f(x^{\star})\leq e^{-2\mu t}(f(x(0))-f(x^{\star})),$
where $x^{\star}$ is the global optimal solution of (1).
Proof.
We denote $P_{A}:=A^{\top}(AA^{\top})^{-1}A,P_{A}^{\perp}=I-P_{A}$. Then the key argument is that according to strong convexity we have:
$f(x)-f(x^{\star})$ $x2\displaystyle\leq\nabla f(x)^{\top}(x-x^{\star})-\frac{\mu}$$x^{\star}||^{2}$
[…]
$\leq\frac{1}{2\mu}\nabla f(x)P_{A}^{\perp}\nabla f(x)+\nabla f(x)P_{A}(x-x^{\star}).$
Further,
$\frac{df}{dt}$ $=-\nabla f(x)^{\top}(\nabla f(x)+J_{h}(x)^{\top}\lambda)$
$=-\nabla f(x)^{\top}(\nabla f(x)-A^{\top}(AA^{\top})^{-1}(A\nabla f(x)-(Ax+b)))\qquad\textup{(}J_{h}(x)=A\textup{)}$
$=-\nabla f(x)^{\top}P_{A}^{\perp}\nabla f(x)-\nabla f(x)^{\top}A^{\top}(AA^{\top})^{-1}A^{\top}(Ax+b)$
$=-\nabla f(x)^{\top}P_{A}^{\perp}\nabla f(x)-\nabla f(x)^{\top}A^{\top}(AA^{\top})^{-1}A^{\top}(Ax-Ax^{\star})$
$=-(\nabla f(x)P_{A}^{\perp}\nabla f(x)+\nabla f(x)P_{A}(x-x^{\star})),$
and thus we have
$\frac{d(f(x(t))-f(x^{\star}))}{dt}\leq-2\mu(f(x(t))-f(x^{\star})),$
which completes the proof ∎
Appendix D Proof of Theorem 2 and 3
@@
is the dual problem of (4) (see Lemma 2 in Appendix G). Thus the solution of $\lambda^{\star}$ in (14) serves as the Lagrangian dual variable of (3.2). Then, we have that the $x^{\star}$ is an optimal solution of (4) if and only if
@@
Appendix E Proof of Theorem 4
@@
Appendix F Proof of Theorem 5
@@
F.1 Intuition: optimization with affine constraints
@@ Appendix G Auxiliaries
$\operatorname*{arg\,min}_{x}\nabla f(x_{t})^{\top}(x-x_{t})+\frac{1}{2}(x-x_{t})^{\top}T(x_{t})^{-1}(x-x_{t})+\lambda^{\top}\left(h(x_{t})+J_{h}(x_{t})(x-x_{t})\right)$
$=x_{t}-\eta T(x_{t})(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda),$
@@
Assumption 3 along with Assumption 1 is a sufficient condition of Assumption 4.
@@
From Assumption 3 and Assumption 1 we have that
@@
Hence, Assumption 4 is satisfied by setting $L=\frac{\lambda_{\max}M^{2}+\|\lambda_{\max}(K)h(x(0))\|}{{\lambda_{\min}}}D^{2}$. ∎